I have worked this problem several times and come up mwith different answers i think i am missing some steps please if someone can work the problem so I can see how you come up with your answers.

Two airplanes are leaving O'Hare at the same time, one bound for New York and the other for California. The one heading to California has a tail wind and is thus going 25 MPH faster than the one headed for New York. If after one hour they are 575 miles apart, what are their speeds?
Eric

Answer by checkley77(12844)

Assuming one is going due West & the other is going due East then their distances are additive.
d=rt
575=(x+25)*1+x*1
575=x+25+x
2x=575-25
2x=550
x=550/2
x=275 mph for the plane headed for New York.
175+25=300 mph for the plane headed for California.
Proof:
275+300=575
575=575